### v2

# ASKAP and VLASS search for a radio-continuum counterpart of ultra-high-energy neutrino event KM3–230213A

Abstract
We present the results of an Australian Square Kilometre Array Pathfinder (ASKAP) 944 MHz and Very Large Array Sky Survey (VLASS) 3 GHz search for a radio-continuum counterpart of the recent ultra-high-energy (UHE) neutrino event, KM3–230213A. Using ASKAP, we catalog 1052 radio sources within the 15 radius search area (68% certainty region) around the particle’s calculated origin, 10 of which we classify as blazar candidates based on their radio spectra. The most prominent radio source in the search area is the nearby spiral galaxy UGCA 127 (nicknamed Phaedra111From Greek: , a Cretan princess of Greek Mythology, derived from Phaidros, Greek: , meaning ’bright’.). Its non-thermal radio spectrum classifies it as a non-blazar active galactic nucleus (AGN). We also present an extended radio source, WISEA J061715.89075455.4 (nicknamed Hebe222From Greek: H, the Greek goddess of youth.), located only 7′ from the geometric center of the search area, with a very unusual highly polarized compact component. Finally, we present a strong radio source, EMU J062248072246 (nicknamed Narcissus333From Greek was a self-absorbed hunter from Thespiae in Boeotia.), which can be modeled either with a synchrotron self-absorbed spectrum or synchrotron emission from thermal electrons. It exhibits flux density variability over the 5-year VLASS 3 GHz survey.
- AGN
- active galactic nucleus
- ASKAP
- Australian Square Kilometre Array Pathfinder
- CASDA
- CSIRO ASKAP Science Data Archive
- CMB
- cosmic microwave background
- CR
- cosmic ray
- CSIRO
- Australian Commonwealth Scientific and Industrial Research Organisation
- EMU
- Evolutionary Map of the Universe
- Gaia DR3
- Gaia Data Release 3
- GLEAM-X
- GaLactic and Extragalactic All-sky MWA eXtended
- GLEAM
- GaLactic and Extragalactic All-sky MWA
- GPS
- gigahertz peaked spectrum
- GRB
- gamma-ray burst
- LIRG
- luminous IR galaxy
- NVSS
- NRAO VLA Sky Survey
- PMN
- Parkes-MIT-NRAO
- POSSUM
- Polarization Sky Survey of the Universe’s Magnetism
- RACS
- Rapid ASKAP Continuum Survey
- RM
- Rotation Measure
- RMS
- root mean squared
- RNe
- Reflection Nebula
- SFR
- star formation rate
- SN
- supernova
- SNR
- supernova remnant
- SMBH
- super massive black hole
- UHE
- ultra-high-energy
- VAST
- Variable and Slow Transients
- VLASS
- Very Large Array Sky Survey
- VLBA
- Very Long Baseline Array
- VLBI
- Very Long Baseline Interferometry
- XRB
- X-ray binary
1 Introduction
Currently, only three astrophysical sources are confirmed as neutrino emitters: our Sun, the supernova SN 1987A, and the Galactic plane (Filipović & Tothill, 2021; Icecube Collaboration et al., 2023). Other very likely neutrino emitters are the blazar TXS 0506+056 and the active galaxy NGC 1068 (IceCube Collaboration et al., 2022). Other possible neutrino origins include cosmogenic, Galactic, and extragalactic sources. Astrophysical neutrinos above GeV energies are typically produced by high-energy cosmic rays interacting with matter or lower-energy photons, and these CRs originate from a variety of sources such as supernova remnants, supernovas, microquasars, star-forming galaxies, gamma-ray bursts, super massive black holes, and AGNs (Filipović & Tothill, 2021). The observed and theorized neutrino energies differ between these sources, with SNs at the lower energies (100 MeV; Hirata et al., 1987), followed by star-forming galaxies (100 GeV up to PeV; Bykov et al., 2015; Fang et al., 2020), SNRs (up to PeV or higher when embedded in dense star clusters; Bykov et al., 2015; Vieu & Reville, 2023), extending up to AGNs (up to EeV; Rodrigues et al., 2021).
The recent KM3Net discovery on 13th February 2025 of the UHE neutrino (KM3–230213A, hereafter KM3–230213A; The KM3NeT Collaboration et al., 2025) provoked a worldwide search for its origin. The UHE neutrino’s origin was pinpointed to within 3° of RA(J2000) = 943, Dec(J2000) = 78 with a 99% certainty, and to within 15 with a 68% certainty. Being close to the equator, it is well-positioned for analysis by numerous radio surveys, including the Evolutionary Map of the Universe (EMU; Norris et al., 2021, Hopkins et al., submitted) and the Polarization Sky Survey of the Universe’s Magnetism (POSSUM; Gaensler et al., 2010, Gaensler et al., submitted) surveys, both conducted with ASKAP (Hotan et al., 2021), as well as the multi-epoch VLASS (Lacy et al., 2020) survey.
Several origins for KM3–230213A have been investigated, including Galactic, which was deemed unlikely (Adriani et al., 2025), a GRB origin (Amelino-Camelia et al., 2025), a cosmogenic origin (The KM3NeT collaboration, 2025; KM3NeT Collaboration et al., 2025), and a blazar origin, specifically the well-known blazar, PMN J0606–0724, which experienced a radio flare within five days of the KM3–230213A event (KM3NeT Collaboration et al., 2025). We here present the results of a search for a potential radio-continuum counterpart of the KM3–230213A event using the ASKAP- Evolutionary Map of the Universe (EMU) survey and the Quick-Look images from VLASS (Lacy et al., 2020). The data used in this paper are explained in Appendix A.
2 Results and Discussion
2.1 Catalog generation
We use the source finding software AEGEAN (Hancock et al., 2018) to catalog ASKAP-EMU radio sources above 10 within the 68% certainty search area. The available EMU data only covered approximately half of the 99% certainty region, and so the 68% certainty region was used. Note that part of the north-western region is not covered (Figure 1). We found 1045 radio point sources, and seven extended sources at 15″ resolution, which we manually measured by fitting regions using the CARTA imaging software, following the methodology described in Filipović et al. (2022, 2024); Smeaton et al. (2024); Bradley et al. (2025). We estimate a 10% error for all ASKAP flux density measurements (Hopkins et al., submitted.).
We use the EMU catalog as a base to cross-match with other catalogs using a 2″ matching radius. We calculate the spectral indices , defined as (Filipović & Tothill, 2021), for all sources with more than two flux density measurements, and cross-match with redshift surveys to find distances. We calculate spectral indices for 185 sources and find cataloged redshifts for one point source (Jones et al., 2009) and three extended sources (Koribalski et al., 2004; Jones et al., 2009; Bilicki et al., 2014). A table of the documented information is provided in Appendix B.
2.2 Objects of interest
We find several objects of interest. The potential blazar counterpart, PMN J0606–0724, analyzed by KM3NeT Collaboration et al. (2025), is not discussed here as it is located outside of the EMU tile coverage.
2.2.1 Radio spectra candidates
We calculated spectral indices for 185 point sources with a mean value of = 0.750.12 (SD = 0.390.15). This value and distribution match the observed average for extragalactic radio point sources (Pennock et al., 2021; Filipović et al., 2021), indicating that the sources have predominantly extragalactic origin.
Blazars are one of the expected astrophysical high-energy neutrino sources, and are typically radio-continuum variable with flat spectral indices (Filipović & Tothill, 2021). We find ten blazar candidates based on their spectra, seven that exhibit a gigahertz peaked spectrum (GPS)-like spectrum (Collier et al., 2018) (EMU J061335–080331, EMU J061750–083920, EMU J061831–075415, EMU J061835–081043, EMU J061848–074327, EMU J062215–075522, and EMU J062245–072330 (see Sec. 2.2.4)), and three that exhibit possible spectral variability (EMU J061227–080810, EMU J061415–083920, and EMU J062244–072334). These blazar candidates are also possible producers of the KM3–230213A event.
2.2.2 Phaedra (UGCA 127; EMU J062056082949)
The most prominent radio source in the field is the nearby Scd type galaxy UGCA 127 (HIPASS J0620–08, see Appendix C), with a systemic velocity of km s-1 (Koribalski et al., 2004) and a distance of Mpc. We give it the nickname Phaedra.
We measure a radio-continuum extent of 210″90″ (8.43.6 kpc) in the ASKAP 944 MHz image (Figure 2, top-right) and detect two resolved hotspot-like features in the 1.3 GHz MeerKAT image. These indicate potential AGN jets oriented within the galaxy plane (marked East and West in Figure 2, top-left). We detect linear polarization in the Polarization Sky Survey of the Universe’s Magnetism (POSSUM) data. Unfortunately, this area is contaminated with bright, smooth Galactic foreground emission at overlapping Rotation Measure (RM)s, meaning we cannot easily separate the features, and so this polarization data is not used in this analysis. We combine our cataloged ASKAP flux densities (Appendix B) with Parkes-MIT-NRAO (PMN) (Griffith et al., 1995, = 0.0840.011 Jy) and MeerKAT (Condon et al., 2021, = 0.2010.020 Jy; B.S. = 7.68″7.52), and calculate an 0.610.01 and a radio luminosity = 1.91021 W Hz-1. This indicates non-thermal emission, consistent with the presence of possible CR accelerators within Phaedra.
We generate a spectral index map with high-resolution EMU, MeerKAT, and 3-epoch averaged VLASS data using the miriad (Sault et al., 1995) maths function. We convolve all images to a common beam size of 10″10″ and calculate the spectral index per pixel. The VLASS image is only reliable for the inner emission due to missing short spacing, so the images were cut at a 2 mJy beam-1 level. We measure an average over the entire region shown (Figure 2, middle left), with the west component being the flattest . This is consistent with a central engine of higher energy injection and particle acceleration in the west component.
We calculate Phaedra’s IR colors using AllWISE data (Cutri et al., 2013) as W1W2 = 0.226 and W2W3 = 3.839. Comparing with the color-color diagram of Wright et al. (2010, their Figure 12), we find Phaedra located in the region common to spiral galaxies, starburst galaxies, and luminous IR galaxies. Being in the starburst region indicates that Phaedra may be experiencing a starburst, potentially caused by AGN feedback, which provides another potential site of particle acceleration (Aiello et al., 2024; Marinelli et al., 2021). While we lack direct observational evidence for star-forming galaxies as UHE neutrino sources (Bechtol et al., 2017), there are numerous models which predict these areas can produce neutrinos up to PeV energies (Bykov et al., 2015).
We detect an offset between the central radio hotspots, IR and optical images, suggesting possible inner knots of a radio jet rather than the galaxy nucleus (Velović et al., 2022). If the emission is caused by jets, it may indicate areas of higher energy if the jets are oriented within the galaxy plane and thus pushing through denser material. It also suggests that the emission may instead be similar to a superbubble, (e.g. 30-Doradus C in the Large Magellanic Cloud; Kavanagh et al., 2015; Sano et al., 2017; Kavanagh et al., 2019; Yamane et al., 2021), microquasars (e.g. Galactic SS 433 or NGC 7793’s S26; Soria et al., 2010; H. E. S. S. Collaboration et al., 2024), recent core-collapse SN event(s) (Aiello et al., 2022; Ando et al., 2023), or massive star clusters (e.g. Westerlund 1; Haubner et al., 2025). All of these are areas of potential particle acceleration and CR or neutrino origin. The likely presence of dense material within the central galaxy provides a natural target for neutrino production by the CRs.
We also suggest the X-ray binary (XRB) 2SXPS J062057.7–082939 (see Figure 1 top-left; Inoue et al., 2021; Walton et al., 2022) located 25″ from Phaedra’s center as a possible origin for KM3–230213A. This source is detected in 2SXP with an X-ray luminosity erg s-1 (2–10 keV range), but is not detected in eRASS1 (Merloni et al., 2024), indicating that it is likely highly X-ray variable, as expected from an active XRB. We see a possible radio counterpart in the MeerKAT image and estimate a radio luminosity 1019 W Hz-1. This would make the object superluminous for an XRB radio counterpart, indicating it may be an area of UHE particle acceleration (Rhodes et al., 2022). Using the ASKAP, MeerKAT and VLASS 3-epoch combined image, we estimate a spectral index, , and detect no observable variability between the three VLASS epochs. We also note three XRB candidates in Phaedra’s outer field (2SXPS J062053.1–083026, 1eRASS J062054.9–083006, 1eRASS J062047.1–083026) that have no corresponding radio emission.
Phaedra is unlikely to contain a blazar due to its steeper spectral index and lack of detected variability. However, it hosts several potential areas of UHE particle acceleration. Due to particle escape as in the commonly used “leaky box” models, the spectrum of the diffuse CRs within spiral galaxies is likely different than the CR spectra injected from the accelerators. Furthermore, reacceleration of CRs, such as for example in the so-called “espresso” model, where the CRs produced at other locations of the galaxy are reaccelerated within the jets of a central AGN engine, can also change the resulting CR distribution and contribute to the production of UHECRs (Caprioli, 2015). These processes might be identified by differentiating areas of different spectral indices within a galaxy, such as done by Duric et al. (1995); Duric (2000). However, our radio data is unfortunately not of sufficiently high resolution for such a detailed spectral analysis. The proposed AGN jets are a potential origin, but they appear oriented away from us, and the star formation rate (SFR) is inconclusive. These jet orientations have been observed in other galaxies, some of which are possible high-energy neutrino sources, e.g. NGC 1068 (Padovani et al., 2024), NGC 7469 (Sommani et al., 2024), and NGC 5938 (Zakir et al. in prep). It is possible that these jets are emitting UHE-CRs into surrounding space; by interacting with ambient gas or photons, they can produce neutrinos. This raises the possibility that the neutrino direction may not point directly to the particle acceleration site.
To assess the probability of chance association of Phaedra, we used the bright galaxy number counts (Yasuda et al., 2001), which implies 0.4 galaxies with a similar or brighter magnitude in a 15 radius. This can mean a likely chance association, but we note that galaxies with radio properties similar to that of Phaedra are rarer than the total number counts.
2.2.3 Hebe (WISEA J061715.89075455.4)
The closest extended radio source to the target position is found 7′ to its south. Called Hebe here, it is a double radio galaxy with diffuse lobes and an extraordinary slightly resolved polarized component, offset by 3″ along the direction of its major axis from the IR galaxy WISEA J061715.89075455.4 (see Figure 3). At a redshift of = 0.1250.017 ( = 553.8938.77 pc; Bilicki et al., 2014), this offset corresponds to 8 kpc. No compact component is visible in the total intensity image, but we made a very rough estimate of its contribution by forcing a fit to a Gaussian component at that position, along with additional components for the lobes. This yielded a flux of 1.5 mJy/(18″ beam).
The bright peak polarized intensity is 0.8 mJy beam-1 above its local background. There is clear depolarization, consistent, e.g., with a Burn-slab model with an rms scatter of . After correction for depolarization, this corresponds to an apparent remarkable 100% polarization, varying only slightly with the unknown spectral index. Even core-dominated quasars have typical fractional polarizations of only a few percent (Pushkarev et al., 2023), although the resolved components of pc-scale Very Long Baseline Interferometry (VLBI) jets can reach levels of 10s of percent (e.g., Jorstad et al., 2005). The rotation measure of the polarized peak is , close to the value of the local Galactic foreground, (; Hutschenreuter et al., 2022), so there is no evidence for a large, additional Faraday component local to Hebe.
At 8 kpc from the presumed AGN core in WISEA J061715.89075455.4, Hebe’s central polarized component could also be a jet, which is only slightly resolved here, although its direction would be almost perpendicular to the more extended emission. We estimate an integrated spectral index for Hebe (see Table 1), although we are not able to measure any spatial-spectral structure, especially of the polarized core.
There were four ASKAP observations of the field spanning December 11th, 2023 through April 13th, 2024, three of which were rejected because of instrumental problems. Nonetheless, we looked for variability in the peak polarized intensity and found it to be . In total intensity, the quality of the data prevented any useful limits.
WISEA J061715.89075455.4 is the northernmost member of a triplet of galaxies, which can be seen in PanSTARRS to be embedded in a common faint envelope. The envelope’s southern boundary has a curious bow shape, perhaps indicating dynamical interactions within this system.
It is unclear whether Hebe is related to the class of blazars or other core-dominated radio AGN. Blazars are of interest as possible sources of cosmic neutrinos (Albert et al., 2024), so further investigation of Hebe is warranted. One interesting model to explore for Hebe would be the “espresso” scheme (Caprioli, 2015) using relativistic jets for the generation of UHECRs. Monitoring in radio polarization and searching for possible optical polarization in WISEA J061715.89075455.4 would be especially valuable.
2.2.4 VLASS investigation and Narcissus (EMU J062248072246)
We search for VLASS counterparts for all cataloged ASKAP sources. We then average the images from the three VLASS epochs and convolve them to the ASKAP resolution of 15″. We record the peak flux densities as VLASS is insensitive to extended emission. We measure the spectral index between ASKAP and VLASS and find for the brightest ASKAP sources, slightly steeper than expected (Pennock et al., 2021; Cotton et al., 2024) likely due to missing VLASS flux densities, and flatter and inverted spectral indices for dimmer ASKAP sources ( 5 mJy) due to our 3 VLASS cutoff. We find six sources with mJy and (inverted), marked with asterisks in Appendix B, which deserve further scrutiny.
We examine the sources with the largest variability in the three VLASS epochs. The median differences between the full-resolution epoch pairs for this VLASS tile can be modeled as a noise contribution of 0.2 mJy beam-1, and a flux normalization error of 5%, arising from instrumental effects including incomplete cleaning. We visually examine 20 sources showing the highest fractional variability integrating the flux density in a 20″ box. We investigate a number of sources which show significant variability.
The most obviously variable source across the VLASS epochs (2017 November 30, 2020 October 9, and 2023 March 10) is the southern source in Figure 4, (EMU J062248072246, hereinafter Narcissus). At the native VLASS 3″ resolution, its peak flux densities are 85.4, 70.4, and 65.2 mJy beam-1 with errors of 0.15 mJy beam-1. The northern source (EMU J062248072246) also appears variable, with peak flux densities of 2.1, 3.7 and 1.3 mJy beam-1. Both components are consistent with being unresolved within the errors. We also search within the EMU data, but detect no variability for Narcissus during the 5-hr observation. Both sources have WISE counterparts (WISEA J062244.90–072334.9 – south, and WISEA J062248.41–072247.9 – north), with colors consistent with AGNs, as expected for radio-selected samples (e.g., Wong et al., 2025). To better separate the adjacent bright WISE source in the south, we looked at the unWISE colors, which yield W1–W2=1.3, consistent with an AGN. Approximately 4% of compact sources at 3 GHz exhibit variability at the 30% level (Mooley et al., 2016), and are expected to be associated with AGN, so these results themselves are not unusual.
What is unusual is the broadband spectra of Narcissus, when we combine 4.85 GHz data (Griffith et al., 1995) with our catalog data (Figure 4, left). We show two possible models. The first model, with the maximum possible low-frequency slope of , corresponds to a compact, homogeneous source with synchrotron self-absorption. At the high, optically thin part of the spectrum, where the variability is observed, the slope is =–0.5, consistent with the limit for strong (non-relativistic) Fermi I shock acceleration (Filipović & Tothill, 2021). Although an AGN connection for Narcissus is attractive, we cannot rule out a galactic origin. We therefore show an alternative model in Figure 4, with a low-frequency slope of . This is similar to shock models explored by Margalit & Quataert (2024), e.g., their Fig. 3(b), used to describe explosive events such as SN1998bw and AT2018cow. Thermal absorption is another possibility for the extremely low-frequency slope.
Independent of the origins of Narcissus, Galactic or extragalactic, the steep low-frequency slope indicates that the emission is dominated by a compact, homogeneous component. This is built-in to explosive models relying on a single shock, whereas inhomogeneities, or multiple components, would broaden the spectrum and flatten the slope. However, it is not clear how the need for a compact, homogeneous source applies to Narcissussince the broadband spectra span many years of observations. Further multi-frequency monitoring and VLBI observations are necessary to separate the spectra of the short- and long-term components and to compare, e.g., to the evolutionary scenarios, e.g., of Margalit & Quataert (2021).
The third VLASS epoch was observed less than one month after the KM3–230213A event, so we conduct an additional search for variability between the VLASS epochs, independent of the ASKAP base catalog, based on an approach similar to Hajela et al. (2019) and Ross et al. (2021). We identify two variable sources with and . The first source is the rotating variable radio star 1RXS J061542.9072339444https://radiostars.org/ (Driessen et al., 2024) which brightens from flux densities of and mJy in the first two epochs, to mJy in the third epoch. The second source is the AGN, WISEA J062248.41072247.9, which has flux densities of 2.10.3, 3.70.4 and 0.2 mJy, in the three VLASS epochs, respectively. Its radio variability and the quasar-like IR colors are consistent with blazars, which are known neutrino emitters (IceCube Collaboration et al., 2018).
2.2.5 VAST investigation
The KM3–230213A event is covered by the ASKAP Variables and Slow Transients (VAST) extragalactic survey (Murphy et al., 2021). The survey began in June 2023 and covers deg2 at a central frequency of 888 MHz to a noise level of Jy approximately every two months.
We queried the VAST transient detection pipeline (Stewart et al., 2024b) for sources within the localization region using vast-tools (Stewart et al., 2024a). We applied standard quality cuts to remove artifacts, requiring the nearest neighbor to be separated by ; a peak-integrated flux ratio of and at least one detection with an S/N greater than seven. 44 sources satisfy these criteria, but after manually inspecting all light curves, none exhibit sufficient variability to be associated with the neutrino event.
Separately, we have searched for variability from the two EMU sources (shown in Figure 4) of interest reported in this work. EMU J062248–072246 is not detected in our survey, which is not surprising given the relative sensitivity. EMU J062244–072333 is detected in all 10 observations which cover it, with a typical integrated flux density of mJy. It exhibits minimal variability, with a modulation index of 8% across the observation, and no obvious trends, with any intrinsic variability inseparable from stochastic variability induced by refractive interstellar scintillation.
3 Conclusion
We catalog ASKAP and VLASS radio sources in the 68% confidence region of the UHE neutrino event, KM3–230213A. We identify ten possible blazar candidates in the field that could produce UHE neutrinos. We further analyze three individual objects, the nearby spiral galaxy UGCA 127 (Phaedra), the radio galaxy (Hebe) closest to the geometric search area center and an unusual compact radio source (Narcissus). Phaedra is the most prominent radio source in the field and is a star-forming galaxy. However, jets may be interacting with the galaxy plane itself, causing starburst regions. This possibility, as well as that of SNRs and microquasars within the galaxy, and the well-known XRB, provide areas of possible neutrino generation within Phaedra. Hebe contains an unusual, very highly polarized compact component, likely associated with one of a triplet of galaxies in a common envelope. Narcissus is a radio source that displays radio variability in VLASS, originating in a compact, heavily absorbed AGN or a more local explosive event. These properties are consistent with known neutrino produces, making Narcissus a plausible candidate for the neutrino origin. Both Hebe and Narcissus are clearly associated with an explosive activity that could produce UHE neutrinos, making them the most likely candidates. Phaedra shows evidence of more high-energy emission than other galaxies, making it a possible candidate, albeit this evidence is somewhat weaker than for the others. All of these systems deserve further investigation as possible origins for KM3–230213A.
4 Acknowledgments
This scientific work uses data obtained from Inyarrimanha Ilgari Bundara / the Murchison Radio-astronomy Observatory. We acknowledge the Wajarri Yamaji People as the Traditional Owners and native title holders of the Observatory site. The Australian Commonwealth Scientific and Industrial Research Organisation (CSIRO)’s ASKAP radio telescope is part of the Australia Telescope National Facility555https://ror.org/05qajvd42. Operation of ASKAP is funded by the Australian Government with support from the National Collaborative Research Infrastructure Strategy. ASKAP uses the resources of the Pawsey Supercomputing Research Centre. Establishment of ASKAP, Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory and the Pawsey Supercomputing Research Centre are initiatives of the Australian Government, with support from the Government of Western Australia and the Science and Industry Endowment Fund. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation. We thank C.W. James for discussion relating to the origin of the neutrino event and Yuri Y. Kovalev for the radio-continuum discussion. This research was funded in part by the National Science Centre, Poland, grant Nos.2018/30/E/ST9/00208 and 2023/49/B/ST9/00066. This research was supported by the Sydney Informatics Hub (SIH), a core research facility at the University of Sydney. This work was also supported by software support resources awarded under the Astronomy Data and Computing Services (ADACS) Merit Allocation Program. ADACS is funded from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government and managed by Astronomy Australia Limited (AAL). We thank an anonymous referee for excellent comments and suggestions, which greatly improved this paper.
5 Data Availability
The ASKAP-EMU data used in this work are available through the CSIRO ASKAP Science Data Archive (CASDA) (https://data.csiro.au/domain/casdaObservation). The full table of ASKAP EMU sources generated for this paper will be made available on VizieR (https://vizier.cds.unistra.fr/viz-bin/VizieR) upon submission.
Appendix A Data
A.1 ASKAP EMU+POSSUM
The 30 deg2 ASKAP tile that includes the KM3–230213A event location was observed as part of EMU (Project Code AS201; Norris et al., 2021, Hopkins et al., submitted) and POSSUM (Project Code AS203, Gaensler et al., 2010, Gaensler et al., submitted). Two ASKAP scheduling blocks (SB 59692; tile EMUA and SB 61077; EMUB) cover this area and were observed on the 2nd of March 2024 and the 13th of April 2024, respectively, i.e. approximately 1 year after the KM3–230213A event. Both tiles use ASKAP’s 36 antenna array with a central frequency of 943.5 MHz, a bandwidth of 288 MHz, and a 300-minute exposure time. Only SB 59692 was used for source finding due to significant interference from a nearby source in SB 61077. However, SB 61077 was still able to be used for polarimetric analysis. The high-resolution EMU image (B.S.=10″) of SB 59692 was also used to generate the spectral index map for UGCA 127 (see Sec. 2.2.2) as the inner components could not be resolved in the convolved 15″ image.
The data were processed using the standard ASKAPSoft pipelines (Guzman et al., 2019), producing both a multifrequency synthesis (MFS) band-averaged Stokes image for EMU, and full Stokes , , and frequency cubes with 1 MHz channels for POSSUM. De-rotation of Stokes and were used to create Faraday spectra, with the amplitude of the peak in the spectrum taken as the polarized intensity. In Figure 1 we show the convolved EMU Stokes radio continuum image which has a resolution of 15″ and an RMS noise sensitivity of = 50 Jy beam-1. We find no Stokes emission within the search area. To obtain Stokes , , polarized intensity (PI) and polarization angle (PA) images, we utilized the RM synthesis technique (Harvey-Smith et al., 2010). We did not use the Fourier transform method from the POSSUM pipeline but the de-rotation technique as described in Ball et al. (2023).
A.2 Other radio catalogs
We use several previous radio catalogs and surveys which overlap the search area to cross-match our sources and include the flux measurements in our catalog to calculate spectral indices (see Table B). We use data from the GaLactic and Extragalactic All-sky MWA eXtended (GLEAM-X) survey (Ross et al., 2024, = 72231 MHz; B.S. 45″), the Rapid ASKAP Continuum Survey (RACS) (Hale et al., 2021, = 888 MHz; B.S. = 25″), the NRAO VLA Sky Survey (NVSS) (Condon et al., 1998, = 1.4 GHz; B.S. = 45″), and the VLASS Quick-Look epoch-averaged images (Gordon et al., 2021, ( = 3 GHz; B.S. 3″)). Limitations for the VLASS Quick-Look images are discussed in Gordon et al. (2021). We also use the different epochs in the VLASS Quick-Look and ASKAP Variable and Slow Transients (VAST) surveys (Murphy et al., 2013, 2021) to search for variability.
Appendix B Example Table
| [1] | [2] | [3] | [4] | [5] | [9] | [10] | [11] | [12] | [13] | [14] | [15] | |
| NAME(S) | RA (J2000) | DEC (J2000) | EMU | GLEAM W087 | GLEAM W215 | RACS | NVSS | VLASS | nα | |||
| 944 (MHz) | 87 (MHz) | 215 (MHz) | 888 (MHz) | 1.40 (GHz) | 3.00 (GHz) | |||||||
| hh:mm:ss.ss | dd:mm:ss.ss | (Jy) | (Jy) | (Jy) | (Jy) | (Jy) | (Jy) | |||||
| aEMU J062056–082949 | 06:20:56.50 | -08:29:49.80 | 0.2550.026 | 1.0170.102 | 0.5590.056 | – | – | – | -0.610.01 | 7 | – | |
| UGCA 127 | ||||||||||||
| Phaedra () | ||||||||||||
| aEMU J061716–075449 | 06:17:16.50 | -07:54:49.60 | 0.0450.005 | 0.2160.022 | 0.1170.012 | – | – | – | -0.650.02 | 7 | 0.130.01 | |
| WISEA J061715.89075455.4 | ||||||||||||
| Hebe (H) | ||||||||||||
| aEMU J062244–072333 | 06:22:44.90 | -07:23:33.40 | 0.1220.012 | 0.019-0.002 | 0.0030.0003 | 0.1100.001 | 0.1060.003 | 0.08280.0046 | 1.090.21 | 8 | – | |
| aEMU J062248–072246 | 06:22:48.40 | -07:22:46.20 | 0.0020.0002 | – | – | – | – | 0.00340.0047 | – | – | – | |
| EMU J062053–084851 | 06:20:53.40 | -08:48:51.70 | 0.1060.011 | – | – | – | – | 0.00240.0007 | – | – | – | |
| EMU J062039–073935 | 06:20:39.90 | -07:39:35.60 | 0.1290.013 | 0.5470.055 | 0.3060.031 | – | – | 0.00790.0013 | -1.100.16 | 7 | 0.130.01 | |
| EMU J061946–064201 | 06:19:46.40 | -06:42:01.10 | 0.0090.001 | 0.1760.018 | 0.0660.007 | – | – | 0.00110.0005 | -1.460.08 | 7 | – | |
| EMU J061944–064232 | 06:19:44.50 | -06:42:32.70 | 0.0110.001 | – | – | – | – | 0.00060.0005 | – | – | – | |
| EMU J061542–063844 | 06:15:42.00 | -06:38:44.50 | 0.1080.011 | 0.5470.055 | 0.2290.023 | – | – | 0.00170.0010 | -1.450.26 | 7 | – | |
| EMU J061931–063426 | 06:19:31.00 | -06:34:26.50 | 0.0020.0002 | – | – | – | 0.0030.0005 | 0.00240.0003 | -0.020.14 | 3 | – | |
| EMU J062118–073729 | 06:21:18.20 | -07:37:29.30 | 0.0070.001 | – | – | – | 0.0040.0004 | 0.00320.0004 | -0.650.29 | 3 | – | |
| EMU J061626–070526 | 06:16:26.80 | -07:05:26.90 | 0.0040.0004 | – | – | 0.0030.0008 | – | 0.00190.0004 | -0.440.32 | 3 | – | |
| EMU J061345–071315 | 06:13:45.10 | -07:13:15.80 | 0.0020.0003 | – | – | 0.0030.0007 | – | 0.00110.0003 | -0.760.11 | 3 | – | |
| EMU J061438–071836 | 06:14:38.60 | -07:18:36.50 | 0.0020.0002 | – | – | 0.0030.001 | – | 0.00110.0004 | -0.640.41 | 3 | – | |
| EMU J061346–082442 | 06:13:46.90 | -08:24:42.80 | 0.0020.0002 | – | – | 0.0030.001 | – | 0.00140.0004 | -0.510.45 | 3 | – | |
| EMU J061356–080149 | 06:13:56.50 | -08:01:49.40 | 0.0040.0004 | – | – | 0.0040.001 | – | 0.00160.0004 | -0.710.10 | 3 | – | |
| EMU J061336–081344 | 06:13:36.90 | -08:13:44.00 | 0.0030.0003 | – | – | 0.0040.001 | – | 0.00190.0003 | -0.420.17 | 3 | – |
Appendix C Hi analysis of the galaxy UGCA 127 (Phaedra)
UGCA 127 (HIPASS J0620–08) shows a symmetric double-horn Hi profile (Figure 5), with a Hi flux density of 250 Jy km s-1 and a Hi mass of . We estimate a rotational velocity of 187 km s-1, suggesting a total dynamical mass of = for an Hi disk radius of 17 kpc (using the Hi mass – size relation, Wang et al., 2016). A gas-rich companion (HIPASS J0622–07) is detected 45′ to the north-east. Inoue et al. (2021) estimate a SFR of 0.36 yr-1 and a stellar mass of .